# Theory of In-Plane-Magnetic-Field-Dependent Excitonic Spectra in Atomically Thin Semiconductors

### v1

@@ Appendix B Parameters
In Tab. 1, we display the parameters used in all analytical (optical spectra) and numerical (Wannier equation) calculations. In particular, we use the momentum-dependent model dielectric function from Ref. [65], which is fitted to ab-initio calculations from the Computational Materials Repository [62], and the screened Coulomb potential derived in Ref. [43] in the Wannier equation for the calculation of the excitonic binding energies and wave functions. To model the screening effect of the h-BN encapsulation, we fit the respective momentum-dependent dielectric function to ab-initio results [66] with static dielectric constant of bulk h-BN $\epsilon_{\text{h-BN},0}=\sqrt{\epsilon_{\text{h-BN},0,\parallel}\epsilon_{\text{h-BN},0,\bot}}=4.8$ [66]. For the light-matter interaction, we use the optical in-plane dielectric constant of h-BN $\epsilon_{\text{h-BN},\infty,\parallel}=4.87$ [67].

### v2

@@ Appendix B Parameters
In Tab. 2, we display the parameters used in all analytical (optical spectra) and numerical (Wannier equation) calculations. In particular, we use the momentum-dependent model dielectric function from Ref. [63], which is fitted to ab-initio calculations from the Computational Materials Repository [2], and the screened Coulomb potential derived in Ref. [11] in the Wannier equation for the calculation of the excitonic binding energies and wave functions. To model the screening effect of the h-BN encapsulation, we fit the respective momentum-dependent dielectric function to ab-initio results [37] with static dielectric constant of bulk h-BN $\epsilon_{\text{h-BN},0}=\sqrt{\epsilon_{\text{h-BN},0,\parallel}\epsilon_{\text{h-BN},0,\bot}}=4.8$ [37]. For the light-matter interaction, we use the optical in-plane dielectric constant of h-BN $\epsilon_{\text{h-BN},\infty,\parallel}=4.87$ [6]. The choice of embedding (h-BN within this manuscript) influences – through its corresponding dielectric constant – the excitonic binding energies, the excitonic wave functions (and therefore the overlaps and Coulomb enhancement) and the radiative linewidth.